# 1in1in1in1in

### v8

@@ 5.2 Open contest at MLCAD 2023
Despite technical differences from infrastructure in [1], the contest offered a great opportunity for Google and the authors of [1] to showcase the versatility and quality of their RL technique. [^12] RL was advertised in [1] as a general technique for combinatorial optimization, so should be able to handle various macro placement formulations. The contest attracted 19 participants, of them 8 finalists — university teams from Taiwan, Hong Kong, China, Canada and the US (several students per team). Google did not participate.131313 Google did not skip a similar research contest at the Int’l Workshop on Logic Synthesis (IWLS 2023). That contest ( https://github.com/alanminko/iwls2023-ls-contest) focused on chip logic design rather than physical design (two nearby fields). Google won the first place using long-running parallel Simulated Annealing. The winning team overlapped with Google Team 1 working on macro placement: Sergio Guadarrama (quoted in [66]), the senior staff software engineer co-authored the ISPD 2022 paper [4] with the lead authors of [1] and then tweeted on 5/3/22 that he and his team at Google “independently replicated” the work in [1].
@@ 6 Rebuttals to critiques of [1]
Despite critical media coverage [53, 54, 60, 64] and technical questions raised, the authors of [1, 3] failed to remove remaining obstacles to reproducibility [63] of the methods and results in [1]. The UCSD team’s engineering effort overcame those obstacles, and they followed up on the work of Google Team 2 [5] that criticized [1], then analyzed many issues listed in Sections 3 and 4. Google Team 2 had access to Google TPU designs and the source code used in [1] before the CT GitHub repo [3] appeared. The UCSD authors of [6, 7] had access to Circuit Training (CT) [3] and benefited from a lengthy involvement of Google Team 1 engineers, but not access to SA code used in [1, 5] or other key pieces of code missing from [3]. Yet, the results in [5] and [6, 7] corroborate each other, and their qualitative conclusions are consistent. UCSD results for Ariane-NG45 closely match those by Google Team 1 engineers, and [7, Figure 4] shows that CT training curves of Ariane-NG45 generated at UCSD match those produced by Google Team 1 engineers. Google Team 1 engineers carefully reviewed the paper [7] and the work in Fall 2022 and Winter 2023, raising no objections [6, FAQ].
The two lead authors of [1] left Google in August 2022, but in March 2023 objected to results of [7] without remedying the deficiencies of [1] (Section 3). Those objections were addressed promptly in [6, FAQ], e.g., in #6, #11, #13, #15. One issue was the lack of pre-training in experiments in [7].
@@
  • Google Team 1 showed [4, Figure 7] that pre-training on “diverse TPU blocks” did not improve results, only runtime. Pre-training on “previous netlist versions” gave small improvement. No such previous versions were discussed, disclosed or released in [1, 3].
@@
Old benchmarks. Another objection [64] is that public circuit benchmarks [22] used in [5, 7] allegedly use outdated infrastructure. In fact, those benchmarks [22] have been evaluated with the HPWL objective, which scales accurately under geometric 2D scaling of chip designs and remains appropriate for all technology nodes (Section 2). Per [2], ICCAD benchmarks were requested [2] by Peer Reviewer #3 of [1]. When [5, 7] implemented this ask, Google RL ran into trouble before routing became relevant: it lost by 20% or so in HPWL optimization (HPWL is the simplest yet most important term of the proxy cost optimized by CT/RL [1, 3]).
@@
Computational resources used in [1] were very large, costly, and difficult to replicate. Since both RL and SA algorithms produce valid solutions early and then gradually improve the proxy function, the best-effort comparisons in [7] used smaller computational resources than in [1], with parity between RL and SA. The result: SA beat RL. [5] compared RL to SA using the same amount of computational resources as in [1], their results were consistent with [5]. If given greater resources, SA and RL are unlikely to improve chip metrics due to poor correlation to the proxy function from [1].
@@ 7 Can the work in [1] be used?
Can Google CT code [3] be improved? RL and SA are orders of magnitude slower than SOTA (Table 3), but pre-training (missing in CT) speeds up RL [4, Figure 7] by only several times.
@@ 8.1 Conclusions about [1]
We crosscheck the results reported in [5, 6, 7] and also account for [2, 3, 4, 50], then summarize conclusions drawn from these works. This confirms many of the initial doubts about [1] and identifies additional deficiencies. As a result, it is clear that [1] is misleading in several ways, such that the readers can have no confidence in the top-line claims of [1]. nor its conclusions. [1] did not improve SOTA while the methods and results of the original paper were not reproducible from the descriptions provided, contrary to stated editorial policies at Nature (see Section 8.3). The reliance on proprietary TPU designs for evaluation, along with ungenerous reporting of experiments, continues to obstruct reproducibility of the methods and the results. The authors of [5] had access to Google internal code whereas [7] reverse-engineered and/or reimplemented missing components. Google Team 2 and the UCSD team drew consistent conclusions from similar experiments, and each team made additional observations.
@@ 8.3 Policy implications
Theoretical arguments and empirical evidence suggest that numerous published papers across various fields do not replicate and are likely false [9, 8, 40, 74]. The developments discussed in this meta-analysis add to the so-called reproducibility crisis that undermines trust in published research results [69, 74]. In response to this crisis, some demand “reproduce or it didn’t happen” [78].
Retraction Watch now tracks 5000 retractions per year, including prominent cases of research misconduct [72, 74]. Per [69], “research misconduct is a serious problem and (probably) getting worse”, which makes it even more important to separate honest mistakes from deliberate exaggerations and misconduct [70, 71, 72, 75]. To this end, see Table 4. Institutional response is needed, and opportunities for future reforms are discussed in [68, 75]. Here we make a modest contribution to this far-reaching discussion by making specific suggestions.
@@
Clearly, when manuscript authors neglect requests for public benchmarking and obstruct reproducibility, their technical claims should be viewed with suspicion [63] (especially if they later disagree with comparisons to their work [59]). Per peer review file [2], the acceptance of the Nature paper was conditional on the release of code and data in the second revision, but this did not happen when the paper was published.
@@ 8.4 Nature editors investigate
Six months later, the investigation has not produced results. And we do not have high hopes, given the Nature News team’s detailed account of an unrelated scandal involving two retracted Nature papers on high-temperature superconductivity by Ranga Dias [79, 80].
@@
Company-internal investigations have even fewer incentives to disclose misconduct to the public.
We believe it is in everyone’s interest to reach clear and unequivocal conclusions about published scientific claims, free of misrepresentations. Authors, Nature editors and reviewers, and the research community, share the burden of responsibility. Seeking the truth is a shared obligation [71, 75]. Separately, Chatterjee is seeking justice [58, 60, 73].

### v9

@@ 5.2 Open contest at MLCAD 2023
Despite technical differences from infrastructure in [1], the contest offered a great opportunity for Google and the authors of [1] to showcase the versatility and quality of their RL technique. [^12] RL was advertised in [1] as a general technique for combinatorial optimization, so should be able to handle various macro placement formulations. The contest attracted 19 participants, of them 8 finalists — university teams from Taiwan, Hong Kong, China, Canada and the US (several students per team). Google did not participate.131313 Google did not skip a similar research contest at the Intl. Workshop on Logic Synthesis (IWLS 2023). That contest ( https://github.com/alanminko/iwls2023-ls-contest) focused on chip logic design rather than physical design (two nearby fields). Google won the first place using long-running parallel Simulated Annealing. The winning team overlapped with Google Team 1 working on macro placement: Sergio Guadarrama (quoted in [67]), the senior staff software engineer co-authored the ISPD 2022 paper [5] with the lead authors of [1] and then tweeted on 5/3/22 that he and his team at Google “independently replicated” the work in [1].
@@ 6 Rebuttals to critiques of [1]
Despite critical media coverage [54, 55, 61, 65] and technical questions raised, the authors of [1, 4] failed to remove remaining obstacles to reproducibility [64] of the methods and results in [1]. The UCSD team’s engineering effort overcame those obstacles, and they followed up on the work of Google Team 2 [6] that criticized [1], then analyzed many issues listed in Sections 3 and 4. Google Team 2 had access to Google TPU designs and the source code used in [1] before the CT GitHub repo [4] appeared. The UCSD authors of [7, 8] had access to Circuit Training (CT) [4] and benefited from a lengthy involvement of Google Team 1 engineers, but not access to SA code used in [1, 6] or other key pieces of code missing from [4]. Yet, the results in [6] and [7, 8] corroborate each other, and their qualitative conclusions are consistent. UCSD results for Ariane-NG45 closely match those by Google Team 1 engineers, and [8, Figure 4] shows that CT training curves of Ariane-NG45 generated at UCSD match those produced by Google Team 1 engineers. Google Team 1 engineers carefully reviewed the paper [8] and the work in Fall 2022 and Winter 2023, raising no objections [7, FAQ].
The two lead authors of [1] left Google in August 2022, but in March 2023 objected to results of [8] without remedying the deficiencies of [1] (Section 3). Those objections were addressed promptly in [7, FAQ], e.g., in #6, #11, #13, #15. One issue was the lack of pre-training in experiments in [8].
@@
  • Google Team 1 showed [5, Figure 7] that pre-training on “diverse TPU blocks” did not improve results, only runtime. Pre-training on “previous netlist versions” gave small improvement. No such previous versions were discussed, disclosed or released in [1, 4].
@@
Old benchmarks. Another objection [65] is that public benchmarks [23] used in [6, 8] allegedly use outdated infrastructure. But those circuits have been evaluated with the HPWL objective, which scales accurately under geometric 2D scaling of chip designs and remains appropriate for all technology nodes (Section 2). Per [2], ICCAD benchmarks were requested [2] by Peer Reviewer #3 of [1]. When [6, 8] implemented this ask, Google RL ran into trouble before routing became relevant: it lost by 20% or so in HPWL optimization (HPWL is the simplest yet most important term of the proxy cost optimized by CT/RL [1, 4]).
@@
Computational resources used in [1] were very large, costly, and difficult to replicate. Since both RL and SA algorithms produce valid solutions early and then gradually improve the proxy function, the best-effort comparisons in [8] used smaller computational resources than in [1], with parity between RL and SA. The result: SA beat RL. [6] compared RL to SA using the same amount of computational resources as in [1]. Their results were consistent with [6]. If given greater resources, SA and RL are unlikely to further improve chip metrics due to poor correlation to the proxy function from [1].
@@ 7 Can the work in [1] be used?
Can Google CT/RL code [4] be improved? RL and SA are orders of magnitude slower than SOTA (Table 3), but pre-training (missing in CT) speeds up RL [5, Figure 7] by only several times.
@@ 8.1 Conclusions about [1]
We crosscheck the results reported in [6, 7, 8] and also account for [2, 4, 5, 51], then summarize conclusions drawn from these works. This confirms many of the initial doubts about [1] and identifies additional deficiencies. As a result, it is clear that [1] is misleading in several ways, such that the readers can have no confidence in the top-line claims of [1]. nor its conclusions. [1] did not improve SOTA while the methods and results of the original paper were not reproducible from the descriptions provided, contrary to stated editorial policies at Nature (see Section 8.3). The reliance on proprietary TPU designs for evaluation, along with insufficient reporting of experiments, continues to obstruct reproducibility of the methods and the results. The authors of [6] had access to Google internal code whereas [8] reverse-engineered and/or reimplemented missing components. Google Team 2 and the UCSD team drew consistent conclusions from similar experiments, and each team made additional observations.
@@ 8.3 Policy implications
Theoretical arguments and empirical evidence suggest that numerous published papers across various fields do not replicate and are likely false [10, 9, 41, 75, 82]. The developments discussed in this meta-analysis add to the so-called reproducibility crisis that undermines trust in published research results [70, 75]. In response to this crisis, some demand “reproduce or it didn’t happen” [79].
Retraction Watch now tracks 5000 retractions per year, including prominent cases of research misconduct [73, 75]. Per [70], “research misconduct is a serious problem and (probably) getting worse”, which makes it even more important to separate honest mistakes from deliberate exaggerations and misconduct [71, 72, 73, 76, 82]. To this end, see Table 4. Institutional response is needed, and opportunities for future reforms are discussed in [69, 76, 83]. Here we make a modest contribution to this far-reaching discussion by making specific suggestions.
@@
Specifically for [1], the Nature editorial [3] insisted that “the technical expertise must be shared widely.” Furthermore, when manuscript authors neglect requests for public benchmarking and obstruct reproducibility, their technical claims should be viewed with suspicion [64] (especially if they later disagree with comparisons to their work [60]). Per peer review file [2], the acceptance of the Nature paper was conditional on the release of code and data in the second revision, but this did not happen when the paper was published or later, per [8]. The Nature paper [1] was amended by the authors to claim that the code had been made available (see the “Data and Code Availability” disclaimer). But serious omissions remain in the released code. This is particularly concerning because ($i$) [1] omitted key comparisons and details, and ($ii$) fraud was alleged under oath in a California court by a Google whistleblower who worked on the project [59]. This makes reproducibility more critical.
@@ 8.4 Nature editors investigate
Nine months later, the investigation has not produced results. And we do not have high hopes, given the Nature News team’s detailed account of an unrelated scandal involving two retracted Nature papers on high-temperature superconductivity by Ranga Dias [80, 81].
@@
Company-internal investigations have even fewer incentives to disclose misconduct to the public, whereas litigation is usually resolved with private settlements and non-disclosure agreements to avoid public admission of wrongdoing.
We believe it is in everyone’s interest to reach clear and unequivocal conclusions about published scientific claims, free of misrepresentations. Authors, Nature editors and reviewers, and the research community, share the burden of responsibility. Seeking the truth is a shared obligation [72, 76].